# Machine learning short-ranged many-body interactions in colloidal systems using descriptors based on Voronoi cells

### v1

@@ II.5 Training and simulation details
To test the result of the trained model, we compare the ML simulations to brute force simulations. These brute-force simulations are MC simulations implemented in the earlier mentioned $N_{C}\mu_{P}AT$ ensemble. All brute-force systems consist of $1024$ colloids and are equilibrated for $10^{7}$ MC steps, after which measurements are collected for $10^{7}$ MC steps.
In order to validate the accuracy of the ML potentials, we studied how well the ML potential reproduced the structure of the system by examining the radial distribution function of the colloid-colloid interactions. Additionally, we measured the free area fractions ($A_{\mathrm{eff}}/A$) to examine whether the potential energies match.
@@ III Results
Figure 8: The radial distribution function $g(r/\sigma_{C})$ as obtained with two different machine learning models that were trained on all input parameters. In the plots, the solid lines represent the ground truth as measured in the system where polymers are treated explicitly (here, the small transparent region indicates twice the standard deviation). In panel a), results are obtained using a model that was trained with the MSLE loss function on a polymer size $\sigma_{P}=0.4$, while in panel b) we used an MSE loss function and trained on a polymer size $\sigma_{P}=0.8$. Both systems have a polymer packing fraction of $\eta_{P}=5.0$ and a colloid packing fraction of $\eta_{C}=0.4$ and both models consists of a 3 hidden layer NN with respectively $[3,3,3]$ nodes. Both the solid lines as well as the plot markers are averaged over five different simulations. In the plot we also indicate the Pearson correlation $\rho$ between true and predicted free area values.
Our aim is to train an ML model that is able to consistently and accurately reproduce the correct structure in a depletion system given a certain polymer size. As mentioned earlier, we test our methodology by comparing the radial distribution functions between the ML model and the full system. Note that reproducing the correct radial distribution function implies that many of the thermodynamic properties are correct hansen2013theory .
@@
In order to test that this combination of network size and loss function leads to models that perform consistently, we trained six independent models with identical hyperparameters for both polymer sizes. As shown in the appendix (right panels of Figs. 11 and 13), this combination of hyperparameters leads to models that are able to consistently reproduce the correct radial distribution function.
To further test the accuracy of our models, for a single model we measured the free area fraction $A_{\mathrm{eff}}/A$ as a function of the polymer packing fraction for the same phase points as displayed in Fig. 6. The results are shown in Fig. 7, where the solid lines represent the full system, while the plot markers represent the ML-systems, and in both cases the results are averaged over five independent simulations. Again, we observe excellent agreement between the full system and the ML potential.
At this point, it is interesting to examine the relative simulation speed obtained from the ML-driven code in comparison to our full-model simulations. Depending on the polymer size and packing fraction, we observe a significant speed up in simulation time between the ML-system and the brute-force system. For example, for $\sigma_{P}=0.4$ and polymer packing fractions above $2.5$ the ML simulations were roughly $10$ to $50$ times faster. This speed-up is all the more impressive when considering the fact that the system we consider here was chosen to be fast to simulate brute-force, while simultaneously having non-trivial many-body contributions.
@@
Loss Functions: As outlined in the methods, we explored the influence of two different loss functions on the accuracy of the models. Between the loss functions that we tested (MSE and MSLE), we found no significant differences (as shown in Appendix Figs. 11 and 13).
Network Size: In contrast to the loss functions, we found that the network size did have an influence on the performance of the models. Although the smaller network size of 3 hidden layers with ($[3,3,3]$) nodes was able to reproduce the correct structure most of the time, we found that its performance was less consistent compared to the larger network size. To explore this, we again trained 6 different models for the smaller network and compared the results to that of the larger network. In Fig. 8 we show the best (green) and worst (red) radial distributions from the six trained models (identified by eye). Here panel a) shows the results for models trained on $\sigma_{C}=0.4\sigma_{C}$ with the MSLE for the phase point $\eta_{P}=5.0$, $\eta_{C}=0.4$, while panel b) shows the results for models trained on $\sigma_{C}=0.8\sigma_{C}$ with the MSE function and $\eta_{P}=5.0$, $\eta_{C}=0.4$ (note that these are the phase points where we saw the models perform most inconsistently). Additionally, in panels c) and d) we repeat this procedure, and pick the best and the worst radial distribution for respectively the same phase points as panels a) en b), but now obtained with the larger NN models. Comparison of the figures clearly indicates the performance of the larger network leads to a significantly more consistent performance. To further demonstrate this, we show results for a wide variety of polymer packing fractions in the Appendix (Figs. 10, 12,11 and 13).
@@ IV Conclusion and outlook
In conclusion, we developed a physically inspired method that is able to consistently predict the effective many-body potential between colloids in a system of hard colloids and ideal polymers. By describing the local environment of each colloid in terms of its associated Voronoi cell and using a simple neural network to predict the potential energy of this cell, we successfully reproduced the structural properties of the system as well as the effective potential energy. Here, we validated the accuracy of our methodology by comparison to brute-force simulations.
@@ Acknowledgements
Data associated with the measurements of the free area fraction and the radial distribution function, as well as all figure data, will be published at the following url [???] prior to publication of the paper.
@@ Appendix A Delaunay- and Voronoi tessellation
Figure 9: a) Cartoon of the representation of a Delaunay tessellation in the Monte Carlo simulation. Note that the red ‘face-centers’ indicate the centers of the circumscribed spheres associated with each triangle.b) Cartoon of a free area prediction. First the Delaunay tessellation (green lines) is converted to a Voronoi tessellation (black lines). Subsequently, we take one of the rhombi consisting of two adjacent triangles that make up the Voronoi cell (b1). We then use the NN to make a prediction for the free area associated with one of the triangles (b2). Finally, the total free area of the rhombus is given by twice the predicted free area (b3).
@@ A.1 Implementing Delaunay tessellation in Monte Carlo simulation
    We set up a linked checklist, which contains all the edges that potentially have to be flipped. Initially, this list contains all the edges of the triangles connected to the displaced particle (e.g. in Fig. 9a, if we would move the middle particle, all depicted edges should be checked). Additionally, we set-up a linked changelist which contains all the edges that are associated with a changed triangle. At the beginning the checklist and changelist are the same.
@@ A.2.1 Periodic boundaries
In our system, we use periodic boundary conditions. This has the advantage that the number of triangles in the DT and thus the number of edges, half-edges and faces stays constant. This does, however, imply that one has to be very careful with which periodic image one uses. Sometimes, in systems with periodic boundaries, one uses nearest-image convention. However, whenever the triangle sides become larger than half the box length, the nearest- image convention no longer holds. To solve this problem, each half-edge carries two numbers which keeps track of which periodic image one should take for that triangle edge, for both the $x$ and $y$ direction. When this number is zero, the half-edge does not cross the boundary of the box, while if it is ±$\pm 1$ it crosses the box in the positive or negative direction. Note that for computing overlap between colloids we use nearest-image convention, as our box is always sufficiently large to allow for this.
@@ Appendix B Different network size and loss function
In this paper, we explored the effect of two different loss functions (MSE and MSLE) and two different network sizes (with 3 and 4 hidden layers respectively) on the accuracy of the ML potential. For each combination of network size and loss function, we trained 6 independent models for two polymer sizes $\sigma_{P}=0.4\sigma_{C}$ and $\sigma_{P}=0.8\sigma_{C}$. In Figs. 10 and 11 we plot the colloid-colloid radial distribution functions for $\sigma_{P}=0.4\sigma_{C}$ obtained with the small and big network respectively. Subsequently, in Figs. 12 and 13 we plot the colloid-colloid radial distribution functions for $\sigma_{P}=0.8\sigma_{C}$ obtained with the small and big network respectively. For all models, we considered a range of different polymer packing fractions. In the figures, panels on the left display results obtained with models trained with the MSE loss function, while the results in the right panels are obtained with models that were trained with the MSLE loss function. In all figures, solid lines represent the radial distribution functions as measured in the full system, where each line is averaged over five independent simulations. The plot markers correspond to the radial distribution function as measured in the ML simulations, where we again averaged over five independent simulations.
@@
Figure 10: The radial distribution functions as obtained with six independently trained models, for $\sigma_{P}=0.4\sigma_{C}$ and respectively an MSE loss function (left panels) and an MSLE loss functions (right panels). Results are obtained with a NN with 3 hidden layers with respectively $[3,3,3]$ nodes. From top to bottom the panels contain results for increasing polymer densities $\eta_{P}$, where $\eta_{P}\in[0.1,2.5,5.0]$. The solid lines represent the ground truth as measured in the system where polymers are treated explicitly (where the transparent region indicates twice the standard deviation), whereas the plotmarkers indicate the radial distribution functions as obtained in the ML system. Both the solid lines as well as the plot markers are averaged over five different simulations. For clarity, different pair correlations are shifted vertically.
@@
Figure 11: The radial distribution functions as obtained with six independently trained models, for $\sigma_{P}=0.4\sigma_{C}$ and respectively an MSE loss function (left panels) and an MSLE loss functions (right panels). Results are obtained with a NN with 4 hidden layers with respectively $[5,5,3,3]$ nodes. From top to bottom the panels contain results for increasing polymer densities $\eta_{P}$, where $\eta_{P}\in[0.1,2.5,5.0]$. The solid lines represent the ground truth as measured in the system where polymers are treated explicitly (where the transparent region indicates twice the standard deviation), whereas the plotmarkers indicate the radial distribution functions as obtained in the ML system. Both the solid lines as well as the plot markers are averaged over five different simulations. For clarity, different pair correlations are shifted vertically.
@@
Figure 12: The radial distribution functions as obtained with six independently trained models, for $\sigma_{P}=0.8\sigma_{C}$ and respectively an MSE loss function (left panels) and an MSLE loss functions (right panels). Results are obtained with a NN with 3 hidden layers with respectively $[3,3,3]$ nodes. From top to bottom the panels contain results for increasing polymer densities $\eta_{P}$, where $\eta_{P}\in[0.1,2.5,5.0,8.0]$. The solid lines represent the ground truth as measured in the system where polymers are treated explicitly (where the transparent region indicates twice the standard deviation), whereas the plotmarkers indicate the radial distribution functions as obtained in the ML system. Both the solid lines as well as the plot markers are averaged over five different simulations. For clarity, different pair correlations are shifted vertically.
@@
Figure 13: The radial distribution functions as obtained with six independently trained models, for $\sigma_{P}=0.8\sigma_{C}$ and respectively an MSE loss function (left panels) and an MSLE loss functions (right panels). Results are obtained with a NN with 4 hidden layers with respectively $[5,5,3,3]$ nodes. From top to bottom the panels contain results for increasing polymer densities $\eta_{P}$, where $\eta_{P}\in[0.1,2.5,5.0,8.0]$. The solid lines represent the ground truth as measured in the system where polymers are treated explicitly (where the transparent region indicates twice the standard deviation), whereas the plotmarkers indicate the radial distribution functions as obtained in the ML system. Both the solid lines as well as the plot markers are averaged over five different simulations. For clarity, different pair correlations are shifted vertically.

### v2

@@ II.5 Training and simulation details
To test the result of the trained model, we compare the ML simulations to brute-force simulations. These brute-force simulations are obtained using both MC and Event Driven Molecular Dynamics (EDMD)rapaport2009event simulations, both implemented in the earlier mentioned $N_{C}\mu_{P}AT$ ensemble, where $N_{C}$ is set to $1024$. The MC simulations are equilibrated for $10^{7}$ MC steps, after which measurements are collected for $10^{7}$ MC steps. The EDMD simulations were run for $t/\tau=10^{5}$ with $\tau=\sqrt{\frac{m\sigma^{2}}{k_{B}T}}$, where one-fourth of this time was used as equilibration time. For the EDMD simulations we adapted the code of Ref. smallenburg2022efficient, .
In order to validate the accuracy of the ML potentials, we studied how well the ML potential reproduces the structure of the system by comparing the radial distribution function of the colloid-colloid interactions. Additionally, we measured the free area fractions ($A_{\mathrm{eff}}/A$) to examine whether the potential energies match. Both of these measurements were tested against the brute-force simulations obtained via MC simulations.
Finally, we compared the virial pressures between the ML system and the brute-force system. The pressure measurement of the brute-force system is performed during an EDMD simulation by tracking the momentum transfer associated with each collisionalder1960studies . In order to measure the pressure in the ML system, we adopted a method analogous to the Widom test insertion methodwidom1963some . Starting from the thermodynamic relation $P=-\frac{\partial F(N,A,T)}{\partial A}$, we approximate it as
$P=-\frac{1}{\Delta A}\left[F(N,A+\Delta A,T)-F(N,A,T)\right],$
where $F(N,A,T)$ is the free-energy for a system with fixed $N$, AA and $T$. By expressing the free energy in terms of the partition function, the free energy difference can be written as
$F(N,$ $\,A+\Delta A,T)-F(N,A,T)$
$=-k_{B}T\log\left(\frac{Z_{A+\Delta A}}{Z_{A}}\right)$
$=-k_{B}T\log\left\langle e^{-\beta\left[U(\mathbf{s}^{N},A+\Delta A)-U(\mathbf{s}^{N},A)-Nk_{B}T\log\frac{A+\Delta A}{A}\right]}\right\rangle,$
where $Z_{A}$ is the partition function of a system with area $A$, where $\mathbf{s}^{N}$ are the scaled colloid coordinates, and where $\langle\dots\rangle$ denotes an ensemble average. Thus, the pressure can be measured during an MC simulation by performing a series of (small) trial area changes and averaging the corresponding potential energy difference. Note that since scaling the system does not affect the Voronoi tesselation, performing a pressure measurement in the ML simulation is computationally cheap.
@@ III Results
Figure 9: Best (green square markers) and worst (red circle markers) radial distribution functions, selected by eye, from six independently models that were trained on all input parameters. In panel a) and c), results are obtained using a model that was trained with the MSLE loss function on a polymer size $\sigma_{P}=0.4$ with respectively 3 hidden layer NN with respectively $[3,3,3]$ nodes and a 5 hidden layer NN with respectively $[5,5,3,3]$ nodes. In panel b) and d), results are obtained using a model that was trained with an MSE loss function on a polymer size $\sigma_{P}=0.8$ the small and large NN size respectively. All systems have a polymer packing fraction of $\eta_{P}=5.0$ and a colloid packing fraction of $\eta_{C}=0.4$. In the plots, the solid lines represent the ground truth as measured in the system where polymers are treated explicitly. Note that in the plot the Pearson correlation $\rho$ between true and predicted free area values is indicated for both runs.
Our aim is to train an ML model that is able to consistently and accurately reproduce the correct structure in a depletion system given a certain polymer size. As mentioned earlier, we test our methodology by comparing the radial distribution functions between the ML model and the full system. Note that reproducing the correct radial distribution function implies that many of the thermodynamic properties are correct hansen2013theory . We also show this thermodynamic consistency explicitly by comparing the virial pressure between the brute-force system and the ML system.
@@
In order to test whether this combination of network size and loss function leads to models that perform consistently, we trained six independent models with identical hyperparameters for both polymer sizes. As shown in the appendix (right panels of Figs. 12 and 14), this combination of hyperparameters leads to models that are able to consistently reproduce the correct radial distribution function.
To further test the accuracy of our models, for a single model we measured the free area fraction $A_{\mathrm{eff}}/A$ as a function of the polymer packing fraction for the same phase points as displayed in Fig. 6. The results are shown in Fig. 7, where the lines represent the full system, the plot markers represent the ML-systems, and where in both cases the results are averaged over five independent simulations. Again, we observe excellent agreement between the full system and the ML potential.
As an additional test of our model, we compare the virial pressures of the ML model and the brute-force model. In Fig. 8, we show the pressure for two different polymer sizes and a range of different colloid and polymer packing fractions (note that we only include polymer and colloid packing fractions that are below colloidal crystallization). As we can see, the two pressure measurements excellently agree with each other.
At this point, it is interesting to examine the relative simulation speed obtained from the ML-driven code in comparison to our full-model simulations. Depending on the polymer size and packing fraction, we observe a significant speed up in simulation time between the ML-system and the brute-force system MC system. For example, for $\sigma_{P}=0.4$ and polymer packing fractions above $2.5$, the ML simulations were roughly $10$ to $50$ times faster. This speed-up is all the more impressive when considering the fact that the system we consider here was chosen to be fast to simulate brute-force, while simultaneously having non-trivial many-body contributions.
@@
Loss Functions: As outlined in the methods, we explored the influence of two different loss functions on the accuracy of the models. Between the loss functions that we tested (MSE and MSLE), we found no significant differences (as shown in Appendix Figs. 12 and 14).
Network Size: In contrast to the loss functions, we found that the network size did have an influence on the performance of the models. Although the smaller network size of 3 hidden layers with ($[3,3,3]$) nodes was able to reproduce the correct structure most of the time, we found that its performance was less consistent compared to the larger network size. To explore this, we again trained 6 different models for the smaller network and compared the results to that of the larger network. In Fig. 9 we show the best (green) and worst (red) radial distributions from the six trained models (identified by eye). Here panel a) shows the results for models trained on $\sigma_{C}=0.4\sigma_{C}$ with the MSLE for the phase point $\eta_{P}=5.0$, $\eta_{C}=0.4$, while panel b) shows the results for models trained on $\sigma_{C}=0.8\sigma_{C}$ with the MSE function and $\eta_{P}=5.0$, $\eta_{C}=0.4$ (note that these are the phase points where we saw the models perform most inconsistently). Additionally, in panels c) and d) we repeat this procedure, and pick the best and the worst radial distribution for respectively the same phase points as panels a) and b), but now obtained with the larger NN models. Comparison of the figures clearly indicates the performance of the larger network leads to a significantly more consistent performance. To further demonstrate this, we show results for a wide variety of polymer packing fractions in the Appendix (Figs. 11, 12, 13 and 14).
@@ IV Conclusion and outlook
In conclusion, we developed a physically inspired method that is able to consistently predict the effective many-body potential between colloids in a system of hard colloids and ideal polymers. By describing the local environment of each colloid in terms of its associated Voronoi cell and using a simple neural network to predict the potential energy of this cell, we successfully reproduced the structural properties of the system as well as the effective potential energy and the virial pressure. Here, we validated the accuracy of our methodology by comparison to brute-force simulations.
@@ Acknowledgements
All simulation codes needed to reproduce the data, as well as all relevant simulation data and notebooks to analyze the data and generate the figures is published at Zenodo prior to publication of the paper: http://doi.org/10.5281/zenodo.15356418 .
@@ Appendix A Delaunay- and Voronoi tessellation
Figure 10: a) Cartoon of the representation of a Delaunay tessellation in the Monte Carlo simulation. Note that the red ‘face-centers’ indicate the centers of the circumscribed circles associated with each triangle. b) Cartoon of a free area prediction. First the Delaunay tessellation (green opaque lines) is converted to a Voronoi tessellation (black lines). Subsequently, we take one of the rhombi consisting of two adjacent triangles that make up the Voronoi cell (b1). We then use the NN to make a prediction for the free area associated with one of the triangles (b2). Finally, the total free area of the rhombus is given by twice the predicted free area (b3).
@@ A.1 Implementing Delaunay tessellation in Monte Carlo simulation
    We set up a linked checklist, which contains all the edges that potentially have to be flipped. Initially, this list contains all the edges of the triangles connected to the displaced particle (e.g. in Fig. 10a, if we would move the middle particle, all depicted edges should be checked). Additionally, we set-up a linked changelist which contains all the edges that are associated with a changed triangle. Note that at the beginning the checklist and changelist are the same.
@@ A.2.1 Periodic boundaries
In our system, we use periodic boundary conditions. This has the advantage that the number of triangles in the DT and thus the number of edges, half-edges and faces stays constant. This does, however, imply that one has to be very careful with which periodic image one uses. Sometimes, in systems with periodic boundaries, one uses nearest-image convention. However, whenever the triangle sides become larger than half the box length, the nearest- image convention no longer holds. To solve this problem, each half-edge carries two numbers which keeps track of which periodic image one should take for that triangle edge, for both the $x$ and $y$ direction. When this number is zero, the half-edge does not cross the boundary of the box, while if it is $\pm 1$ it crosses the box in the positive or negative direction. Note that for computing overlap between colloids we use nearest-image convention, as our box is always sufficiently large to allow for this.
@@ Appendix B Different network size and loss function
In this paper, we explored the effect of two different loss functions (MSE and MSLE) and two different network sizes (with 3 and 4 hidden layers respectively) on the accuracy of the ML potential. For each combination of network size and loss function, we trained 6 independent models for two polymer sizes $\sigma_{P}=0.4\sigma_{C}$ and $\sigma_{P}=0.8\sigma_{C}$. In Figs. 11 and 12 we plot the colloid-colloid radial distribution functions for $\sigma_{P}=0.4\sigma_{C}$ obtained with the small and big network respectively. Subsequently, in Figs. 13 and 14 we plot the colloid-colloid radial distribution functions for $\sigma_{P}=0.8\sigma_{C}$ obtained with the small and big network respectively. For all models, we considered a range of different polymer packing fractions. In the figures, panels on the left display results obtained with models trained with the MSE loss function, while the results in the right panels are obtained with models that were trained with the MSLE loss function. In all figures, solid lines represent the radial distribution functions as measured in the full system, where each line is averaged over five independent simulations. The plot markers correspond to the radial distribution function as measured in the ML simulations, where we again averaged over five independent simulations.
@@
Figure 11: The radial distribution functions as obtained with six independently trained models, for $\sigma_{P}=0.4\sigma_{C}$ and respectively an MSE loss function (left panels) and an MSLE loss functions (right panels). Results are obtained with a NN with 3 hidden layers with respectively $[3,3,3]$ nodes. From top to bottom the panels contain results for increasing polymer densities $\eta_{P}$, where $\eta_{P}\in[0.1,2.5,5.0]$. The solid lines represent the ground truth as measured in the system where polymers are treated explicitly, whereas the plotmarkers indicate the radial distribution functions as obtained in the ML system. Both the solid lines as well as the plot markers are averaged over five different simulations. For clarity, different pair correlations are shifted vertically.
@@
Figure 12: The radial distribution functions as obtained with six independently trained models, for $\sigma_{P}=0.4\sigma_{C}$ and respectively an MSE loss function (left panels) and an MSLE loss functions (right panels). Results are obtained with a NN with 4 hidden layers with respectively $[5,5,3,3]$ nodes. From top to bottom the panels contain results for increasing polymer densities $\eta_{P}$, where $\eta_{P}\in[0.1,2.5,5.0]$. The solid lines represent the ground truth as measured in the system where polymers are treated explicitly, whereas the plotmarkers indicate the radial distribution functions as obtained in the ML system. Both the solid lines as well as the plot markers are averaged over five different simulations. For clarity, different pair correlations are shifted vertically.
@@
Figure 13: The radial distribution functions as obtained with six independently trained models, for $\sigma_{P}=0.8\sigma_{C}$ and respectively an MSE loss function (left panels) and an MSLE loss functions (right panels). Results are obtained with a NN with 3 hidden layers with respectively $[3,3,3]$ nodes. From top to bottom the panels contain results for increasing polymer densities $\eta_{P}$, where $\eta_{P}\in[0.1,2.5,5.0,8.0]$. The solid lines represent the ground truth as measured in the system where polymers are treated explicitly, whereas the plotmarkers indicate the radial distribution functions as obtained in the ML system. Both the solid lines as well as the plot markers are averaged over five different simulations. For clarity, different pair correlations are shifted vertically.
@@
Figure 14: The radial distribution functions as obtained with six independently trained models, for $\sigma_{P}=0.8\sigma_{C}$ and respectively an MSE loss function (left panels) and an MSLE loss functions (right panels). Results are obtained with a NN with 4 hidden layers with respectively $[5,5,3,3]$ nodes. From top to bottom the panels contain results for increasing polymer densities $\eta_{P}$, where $\eta_{P}\in[0.1,2.5,5.0,8.0]$. The solid lines represent the ground truth as measured in the system where polymers are treated explicitly, whereas the plotmarkers indicate the radial distribution functions as obtained in the ML system. Both the solid lines as well as the plot markers are averaged over five different simulations. For clarity, different pair correlations are shifted vertically.